OpenStudy (anonymous):

Which expression is equivalent to (-2)(a+6) -2a+6 2a+12 -2a-12 -2a+12 I think it's the first one, -2a + 6.

OpenStudy (texaschic101):

-2(a + 6) = -2 * a + (-2) * 6 = -2a + (-12) = -2a - 12

OpenStudy (anonymous):

I think it is the second one. Using the Distributive Property, you can multiply -2 and a. You will get -2a. Then you multiply -2 and 6 and you get -12. -2a-12 :)

OpenStudy (anonymous):

it is -2a-12 steps: -2(a + 6) = -2 * a + (-2) * 6 = -2a + (-12) = -2a - 12

OpenStudy (anonymous):

@texaschic101 is correct @wyq12

OpenStudy (anonymous):

So the answer is -2a - 12